OpenStudy (anonymous):

if you save one dollar on january 1,two dollars on january 2,three dollars on january 3, and continue this pattern for one year(not a leap year), what will be the value of your entire savings,in dollars, at the end of that one year?

OpenStudy (anonymous):

This is an arithmetic progression \[a_{n}=a_{1}+(n-1)d=1+(364)1=365\] The sum of a finite arithmetic progression is \[\S_{n}=\frac{n}{2}(a_{1}+a_{n})=\frac{365}{2}(1+365)=$66,795\]